OpenStudy (anonymous):

Emilio has 6 more nickels than dimes, and the total value of his coins is $2.10. How many nickels does he have? Let n represent the number of nickels 0.10n + 0.05n + 6 = 2.10 0.10n + 0.05(n + 6) = 2.10 0.10(n + 6) + 0.05n = 2.10 0.10(n 6) + 0.05n = 2.10

zepdrix (zepdrix):

6 more nickles than dimes, hmm so let's try to write this as an equation. let N be nickels, D be dimes. N + 6 = D <-- Nickles plus 6 more equals the number of dimes. We'll have to make another equation, based on the VALUE of the coins. Dimes are worth 10 cents each. So if we have D dime, we can write it like this, .10D <-- This represents the value of D dimes. Doing the same with nickles gives us, .05N <-- The value of N nickles We'll make an equation out of the TOTAL VALUE of the coins that they gave us at the start. All of the coins should add up to $2.10. N nickles plus D dimes (at their respective values) should add up to 2.10. We can write this as, .05N + .10D = 2.10 So we have a system of equations,\[\huge N+6=D\] \[\huge .05N+.10D=2.10\] Does this much make sense so far?

OpenStudy (anonymous):

these are the options for question 0.10n + 0.05n + 6 = 2.10 0.10n + 0.05(n + 6) = 2.10 0.10(n + 6) + 0.05n = 2.10 0.10(n 6) + 0.05n = 2.10

zepdrix (zepdrix):

Did I do the first one backwards? Maybe it's suppose to be N=D+6, my brain might be thinking a little cooky today :)

OpenStudy (anonymous):

there is no d unfortunately

OpenStudy (anonymous):

ill edit my question

zepdrix (zepdrix):

We need to take this system, and combine them, to give us one equation involving N. I just wanna make sure I didn't make a silly mistake real quick. Yah it should be N=D+6, my bad. Which we can rewrite as N-6=D We want to plug in this value for the D in the other equation. \[\large N-6=(D)\]\[\large .05N+.10(D)=2.10 \qquad \rightarrow \qquad .05N+.10(N-6)=2.10\]

zepdrix (zepdrix):

What is the last option? the symbol isn't showing up

OpenStudy (anonymous):

minus